# Any-Resolution AI-Generated Image Detection by Spectral Learning

### v1

@@ 4.3 Ablation Studies
Pretraining Approach: We evaluate the importance of the frequency reconstruction pretext task for learning the spectral distribution of real images. To this end, we replace our ViT backbone, that has been trained using the process described in Sec. 3.1, and retrain our network using the CLIP [56] and DINOv2 [53] backbones, that have been trained using visual-text alignment and spatial alignment respectively. We report results in Tab. 2 and find that our ViT leads to clearly higher performance, while being trained only on a fraction of the data used for training CLIP and DINOv2.
@@
with chromatic augm. 80.5
Table 3: Ablation studies of the key components. Average AUC over 5 sources of real images and 13 generative models is reported. Best value is highlighted in bold.
@@
Architectural Components: To verify the significance of the main components of our architecture we remove each of them and retrain the network. We report the respective AUC scores in Tab. 3. When removing SRS the performance plunges indicating its significance in detecting AI-generated images as out-of-distribution samples to the underlying spectral model. Removing SCV significantly decreases performance highlighting the significance of spectral context. Last, we remove SCA and evaluate using center cropping of the images as well as ten-crop along with mean and max fusion of the subsequent outputs. In all cases a significant performance drop is visible, highlighting the importance of SCA.
Augmentations: We evaluate our augmentation policy by removing the noise and blur distortions, as well as the JPEG compression augmentation and report performance in Tab. 3. Performance drops in both cases, highlighting their value. To evaluate whether further augmentations could benefit performance, we also train our network using WebP compression and chromatic augmentations. However, we found that this led to a decrease in the overall performance.
Hyperparameter Tuning: We examined different values for the masking radius $r$, the learning rate $lr$ and the size of the latent dimensionality $D$ and we report the corresponding results in Tab. 4. We see that $lr=5e-4$ was optimal for our network, as well as $D=1024$. Regarding the masking radius, the optimality of $r=16$ highlights the importance of the alignment between the pretext task and the later use of this underlying spectral model for AID.
@@ 4.4 Analysis of Spectral Context Attention
We perform a qualitative evaluation of spectral context attention and present two samples in Fig. 4, where we overlay the normalized attention score for each patch of the image. In the first one, depicting an AI-generated image of a human subject possessing a six fingers hand, our attention mechanism identified this patch as significant for the final decision. In the second case, we observe that spectral context attention predicts higher values for patches depicting some complex scenes, in contrast to patches with flat textures aligning without our original intuition for building it.
@@ 4.5 Failure Cases
[Accurately detected AI-generated images shared online.]   [Derivative AI-generated images failed to be detected.]
Figure 5: Failures in detecting derivative AI-generated images.
While our method, based on spectral learning, exhibits superior generalization performance both w.r.t state-of-the-art approaches as well as on robustness against online perturbations, it is still not capable of fully solving the AID task. As Dufour et al. [17] highlighted, AI-generated images circulating online usually appear in the form of derivative images, i.e. parts of synthetic images appearing in screenshots, memes, or even in photographs of computer screens and in printed material. So, the intermediate medium, either fully digital, or even involving analog parts in the case of external photographs and printings, heavily distorts the spectral distribution of the AI-generated image signal. Thus, the subtle inconsistencies introduced by the generative approaches, become indistinguishable in this later noisy signal. Case in point, we present two AI-generated images in Fig. 5 that have been recently shared online. While our method successfully detects the early-shared copies of these images, it fails to detect two of their derivative copies.
@@ 5 Conclusion
In this paper we introduced the key idea that the spectral distribution of real images constitutes an invariant and highly-discriminative pattern for the task of AI-generated image detection and proposed masked spectral learning to model it. To detect generated images as out-of-distributions samples of this model, we introduced the concept of spectral reconstruction similarity for capturing the divergence of the spectral distribution of an image from the one learned by the underlying model. Furthermore, we introduced spectral context attention for detecting subtle spectral inconsistencies in images of any resolution. Using our approach we achieved a significant improvement in state-of-the-art performance, while achieving robustness against common online perturbations. By introducing our work we believe to contribute in reducing the malicious exploitation of generative AI as well as in providing the research community with useful ideas and building blocks for subsequent works in the field.
@@
Acknowledgments: This work was supported by the Horizon Europe vera.ai project (grant no.101070093). The computational resources were granted with the support of GRNET.
@@ 6.1 Projection Operators
$\mathcal{P}_{1}$ and $\mathcal{P}_{2}$: To build the spectral context vector based on the learnable spectral map $C$ we used the projection operators $\mathcal{P}_{1}(\cdot):R^{2D}\rightarrow R^{D}$ and $\mathcal{P}_{1}(\cdot):R^{D}\rightarrow R^{D}$ that share the same architecture. In particular, they first process input features using two linear layers with GELU [31] activations and then normalize their output using a LayerNorm [43] layer.
@@ 6.2 Implementation and Training Details
We implement our approach using PyTorch and train it for 35 epochs using the AdamW [49] optimizer. In the first five epochs we perform a linear warmup of the learning rate ($lr$) and increase it from $72.5e-$ 7 to $45e-$ 4. Then, from the 5th epoch we apply a per-step cosine decay to ultimately decrease it to $72.5e-$ 7 at the last step of the 35th epoch. We set the latent dimensionality $D=1024$ and the masking radius $r=16$ according to the hyperparameter tuning procedure that we presented in Sec. 4.3 of the main paper. The vision transformer (ViT) [16] that we employ in our model of real images $G$ uses a patch size $p=16$ and includes $N=12$ transformer [70] blocks with a latent dimensionality $d=768$. To compute the 2D Discrete Fourier Transform we use the Fast Fourier Transform algorithm.
We split each image into $K$ patches of size $h=w=224$, while we empirically set the latent dimensionality of the spectral context attention to $D_{h}=1536$. During training we employ $K_{training}=4$ patches. These patches are generated as augmented views of the input image, using random resizing, cropping, rotation, Gaussian blur, Gaussian noise and JPEG compression augmentations. The training of spectral reconstruction similarity, spectral context vector and spectral context attention is performed on a single Nvidia L40S 48GB GPU, using half-precision arithmetic, and takes about 50 hours.
@@ 7 Runtime Analysis
To evaluate the computational performance of our approach we analyze its runtime using the proposed spectral context attention as well as by solely relying on the scaled dot product self-attention of the vision transformer [16]. We report the runtime across images of different resolution in Fig. 6. As we see, the runtime increases rapidly when relying on the self-attention of quadratic computational complexity, while, due to memory constraints, the maximum image resolution we could process using an Nvidia L40S 48GB GPU was limited to 18 megapixels. Instead, employing spectral context attention enables our approach to scale linearly w.r.t the size of the image, without being limited by the available memory. To better highlight the difference in memory requirements we present in Tab. 7 the required GPU memory for different image sizes. We see that for a 10 megapixels image self-attention requires almost 18 gigabytes of memory. Instead, when using spectral context attention less than 1 gigabyte of memory is required. Finally, using spectral context attention we managed to analyze with our architecture images of 1 gigapixel, under a single-GPU setup, exceeding by a large margin the size of the images produced by commercial cameras at the time of writing this manuscript. To eliminate any possible inconsistencies due to inefficient implementations, we employed the cuda-optimized FlashAttention [12] throughout our runtime analysis.
@@ 10 Qualitative Evaluation
We perform a qualitative evaluation of our approach across all the considered datasets and present samples for the 13 generative models in Figs. 12, 13, 14, 8, 9, 15, 16, 10, 17, 18, 19, 11 and 20 as well as for the 5 sources of real images in Figs. 21, 22, 23, 24 and 25. As we see, our approach accurately detects images generated by all the considered generative approaches, depicting a diverse set of topics and incorporating different levels of visual fidelity and aesthetics. At the same time SPAI accurately classifies real images originating from all the employed sources. Therefore, employing spectral learning enables our architecture to not rely on some high-level semantics, but to effectively detect the subtle inconsistencies introduced by the generative models, to distinguish between AI-generated and real imagery.

### v2

@@ 4.3 Ablation Studies
Pretraining Approach: We evaluate the importance of the frequency reconstruction pretext task for learning the spectral distribution of real images. To this end, we replace our spectral learned ViT backbone and retrain our network using the CLIP [57] and DINOv2 [54], that have been trained using visual-text alignment and spatial alignment respectively. We report results in Tab. 2 and find that our ViT leads to clearly higher performance, while being trained only on a fraction of the data used for training CLIP and DINOv2.
@@
with chromatic augmentations 80.5
Table 3: Ablation studies of the key components and augmentations. Average AUC over 5 sources of real images and 13 generative models is reported. Best value is highlighted in bold.
@@
Architectural Components: To verify their significance we remove each of them and retrain the network. We report the respective AUC scores in Tab. 3. First, we train the backbone along with the rest of the network from scratch, without using pretrained weights. As expected, performance plunges, highlighting the importance of learning the spectral distribution of real images first. Then, we remove SRS, SCV and SCA. For the latter, we evaluate using center cropping of the images as well as ten-crop along with mean and max fusion of the subsequent outputs. In all cases performance drops, highlighting the importance of these components in detecting AI-generated images as out-of-distribution samples to the underlying spectral model.
Augmentations: We evaluate our augmentation policy by removing the noise and blur distortions, the JPEG compression as well as both and report performance in Tab. 3. Performance drops in all cases, highlighting their value. To evaluate whether further augmentations could benefit performance, we also train our network using WebP compression and chromatic augmentations. However, we found that this led to a decrease in the overall performance.
Hyperparameters Tuning: We examined different values for the masking radius $r$, the learning rate $lr$ and the size of the latent dimensionality $D$ and we report the corresponding results in Tab. 4. We see that $lr=5e-4$ was optimal for our network, as well as $D=1024$. Regarding the masking radius, the optimality of $r=16$ highlights the importance of the alignment between the pretext task and the later use of this underlying spectral model for AID.
@@ 4.4 Analysis of Spectral Context Attention
We perform a qualitative evaluation of spectral context attention and present two samples in Fig. 4, where we overlay the normalized attention score for each patch of the image. In the first one, depicting an AI-generated image of a human subject possessing a six fingers hand, our attention mechanism identified this patch as significant for the final decision. In the second case, we observe that spectral context attention predicts higher values for patches depicting some complex scenes, in contrast to patches with flat textures, aligning with our original intuition for building it.
@@ 4.5 Failure Cases
While our method, based on spectral learning, exhibits superior generalization performance both w.r.t state-of-the-art approaches as well as on robustness against online perturbations, it still cannot claim to fully solve the AID task. As Dufour et al. [17] highlighted, AI-generated images often appear online in the form of derivative images, i.e. parts of synthetic images appearing in screenshots, memes, or even in photographs of computer screens and in printed material. This intermediate medium, either fully digital, or involving analog parts in the case of external photographs and printings, distorts the spectral distribution of the AI-generated image signal. Thus, the subtle inconsistencies introduced by the generative approaches, become indistinguishable in this later noisy signal. Case in point, we present two AI-generated images in Fig. 5 that have been recently shared online. While our method successfully detects their early-shared copies, it fails to detect two of their derivative copies. This highlights some potential future directions on combining spectral learning with semantic context understanding.
[Accurately detected AI-generated images shared online. Spectral context attention correctly identifies artifacts in the hand and the texture of the jacket (left), as well as on the face and the ear (right).]
[Derivative AI-generated images failed to be detected. The distortion introduced by the intermediate medium makes indistinguishable the spectral artifacts that were previously captured correctly.]
Figure 5: Failures in detecting derivative AI-generated images. Similar to Fig. 4, spectral context attention is depicted in overlay.
@@ 5 Conclusion
In this paper we introduced the key idea that the spectral distribution of real images constitutes an invariant and highly-discriminative pattern for AI-generated image detection and proposed masked spectral learning to model it. To detect generated images as out-of-distributions samples of this model, we introduced the concept of spectral reconstruction similarity for capturing the divergence of the spectral distribution of an image from the one learned by the underlying model. Furthermore, we introduced spectral context attention for detecting subtle spectral inconsistencies in images of any resolution. Using our approach we achieved a significant improvement in state-of-the-art performance, while achieving robustness against common online perturbations. By introducing our work we believe to contribute in reducing the malicious exploitation of generative AI as well as in providing the research community with useful ideas and building blocks for subsequent works in the field.
@@
Acknowledgments
This work was partly supported by the Horizon Europe projects ELIAS (grant no. 101120237) and vera.ai (grant no. 101070093). The computational resources were granted with the support of GRNET.
@@ 6.1 Projection Operators
$\mathcal{P}_{1}$ and $\mathcal{P}_{2}$: To build the spectral context vector based on the learnable spectral map $C$ we used the projection operators $\mathcal{P}_{1}(\cdot):R^{2D}\rightarrow R^{D}$ and $\mathcal{P}_{1}(\cdot):R^{D}\rightarrow R^{D}$ that share the same architecture. In particular, they first process input features using two linear layers with GELU [32] activations and then normalize their output using LayerNorm [44].
@@ 6.2 Implementation and Training Details
We implement our approach using PyTorch and train it for 35 epochs using the AdamW [50] optimizer. In the first five epochs we perform a linear warmup of the learning rate ($lr$) and increase it from $72.5e-$ 7 to $45e-$ 4. Then, from the 5th epoch we apply a per-step cosine decay to ultimately decrease it to $72.5e-$ 7 at the last step of the 35th epoch. We set the latent dimensionality $D=1024$ and the masking radius $r=16$ according to the hyperparameter tuning procedure that we presented in Sec. 4.3 of the main paper. The vision transformer (ViT) [16] that we employ in our model of real images $\mathcal{G}$ uses a patch size $p=16$ and includes $N=12$ transformer [71] blocks with a latent dimensionality $d=768$. To compute the 2D Discrete Fourier Transform we use the Fast Fourier Transform algorithm.
We split each image into $K$ patches of size $h=w=224$, while we empirically set the latent dimensionality of the spectral context attention to $D_{h}=1536$. During training we employ $K_{training}=4$ patches. These patches are generated as augmented views of the input image, using random resizing, cropping, rotation, Gaussian blur, Gaussian noise and JPEG compression augmentations. The training of spectral reconstruction similarity, spectral context vector and spectral context attention is performed on a single Nvidia L40S 48GB GPU, using mixed-precision arithmetic, and takes about 50 hours.
@@ 7 Runtime Analysis
To evaluate the computational performance of our approach we analyze its runtime using the proposed spectral context attention as well as by solely relying on the scaled dot product self-attention of the vision transformer [16]. We report the runtime across images of different resolution in Fig. 6. As we see, the runtime increases rapidly when relying on the self-attention of quadratic computational complexity, while, due to memory constraints, the maximum image resolution we could process using an Nvidia L40S 48GB GPU was limited to 18 megapixels. Instead, employing spectral context attention enables our approach to scale linearly w.r.t the size of the image, without being limited by the available memory. To better highlight the difference in memory requirements, we present in Tab. 7 the required GPU memory for different image sizes. We see that for a 10 megapixels image self-attention requires almost 18 gigabytes of memory. Instead, when using spectral context attention less than 1 gigabyte of memory is required. Finally, using spectral context attention we managed to analyze with our architecture images of 1 gigapixel, under a single-GPU setup, exceeding by a large margin the size of the images produced by commercial cameras at the time of writing this manuscript. To eliminate any possible inconsistencies due to inefficient implementations, in our runtime analysis we employed the cuda-optimized FlashAttention [12].
@@
8 Feature Space Analysis
Image Size < 0.5 MPixels 0.5 - 1.0 MPixels > 1.0 MPixels Acc.
Approach Glide SD1.3 SD1.4 Flux DALLE2 SD2 SDXL SD3 GigaGAN MJv5 MJv6.1 DALLE3 Firefly AVG
Dire [73] 38.2 57.7 58.5 46.0 51.5 61.7 47.3 49.6 40.9 43.8 50.4 62.6 51.1 50.7
CNNDet. [72] 52.0 52.2 52.3 49.1 58.0 52.2 56.6 47.5 63.0 50.8 52.6 47.6 55.5 53.0
NPR [67] 79.9 79.9 79.4 29.9 29.9 30.1 29.9 79.9 79.9 31.4 29.9 79.9 29.9 53.1
Fusing [35] 54.1 52.9 53.0 50.5 59.4 52.6 51.4 48.2 62.6 53.2 58.6 48.2 55.9 53.9
FreqDet. [24] 46.9 80.4 80.6 42.6 42.8 46.3 61.8 65.1 60.9 44.0 41.6 46.9 77.2 56.7
UnivFD [53] 50.8 67.9 67.4 45.9 81.0 73.2 65.5 45.0 73.6 50.0 53.8 44.9 88.2 62.1
LGrad [66] 69.3 78.6 78.8 68.9 79.2 54.6 64.3 34.0 80.4 63.3 73.6 37.2 42.1 63.4
GramNet [49] 72.7 78.7 79.1 73.9 80.4 58.5 72.3 32.4 80.4 58.2 80.3 37.0 34.8 64.5
DeFake [64] 76.3 59.5 59.0 79.6 45.2 61.7 51.7 78.3 67.0 62.2 77.1 81.5 44.0 64.8
DMID [7] 52.4 99.3 99.3 82.1 49.5 97.9 96.5 57.9 54.3 98.4 78.1 49.3 57.5 74.8
PatchCr. [79] 69.6 86.6 86.7 79.7 71.7 86.8 89.4 40.2 89.0 72.6 87.9 40.6 72.5 74.9
RINE [40] 88.5 96.6 96.4 84.1 82.0 89.6 95.5 47.0 83.3 90.1 69.2 47.6 67.0 79.8
SPAI (Ours) 81.3 92.2 92.3 71.9 83.4 88.8 90.5 61.3 76.2 87.2 74.3 80.8 90.3 82.3
Table 8: Comparison against state-of-the-art. Average accuracy over 5 sources of real images is reported. Lower values are highlighted in red, while higher values are highlighted in green. Best overall average value is highlighted in bold, while second best is underlined.
Image Size < 0.5 MPixels 0.5 - 1.0 MPixels > 1.0 MPixels AP
Approach Glide SD1.3 SD1.4 Flux DALLE2 SD2 SDXL SD3 GigaGAN MJv5 MJv6.1 DALLE3 Firefly AVG
Dire [73] 39.5 56.6 57.4 45.2 52.2 70.2 47.8 48.0 41.7 44.1 40.1 59.7 48.6 50.1
CNNDet. [72] 58.0 57.6 59.4 44.1 70.8 57.2 67.0 37.8 75.6 50.6 47.2 35.5 69.2 56.1
FreqDet. [24] 46.3 90.6 90.8 42.6 45.2 46.7 60.5 64.7 63.7 43.6 30.5 47.5 71.7 57.3
NPR [67] 75.4 87.6 72.3 39.3 31.3 35.2 38.3 72.4 84.0 38.2 31.0 97.4 50.4 57.9
Fusing [35] 64.7 63.6 63.0 57.7 77.1 65.6 61.9 42.1 80.7 64.6 67.6 36.6 74.0 63.0
LGrad [66] 80.1 76.8 78.4 71.7 83.1 59.3 69.2 32.6 89.7 67.2 69.3 38.3 44.6 66.2
GramNet [49] 74.9 79.4 79.8 74.6 80.3 65.2 74.0 36.2 80.1 63.5 73.3 48.0 45.5 67.3
UnivFD [53] 62.6 82.1 81.9 43.3 91.4 86.1 79.9 38.6 87.0 57.9 52.9 39.6 95.9 69.2
DeFake [64] 87.0 63.0 62.5 90.8 44.9 66.7 54.7 87.4 73.9 66.6 83.8 93.2 42.9 70.6
PatchCr. [79] 79.7 96.3 96.7 86.8 80.0 96.1 95.3 41.5 97.7 81.1 94.1 39.0 77.5 81.7
DMID [7] 71.5 100.0 100.0 97.2 54.9 99.7 99.7 71.8 70.2 99.9 93.2 45.3 87.7 83.9
RINE [40] 95.5 99.9 99.9 94.0 93.2 96.8 99.4 46.9 93.4 97.0 78.0 48.6 83.1 86.6
SPAI (Ours) 90.9 99.3 99.4 82.0 90.7 96.8 97.2 72.4 84.7 95.1 79.1 88.8 94.9 90.1
Table 9: Comparison against state-of-the-art. Average precision over 5 sources of real images is reported. Lower values are highlighted in red, while higher values are highlighted in green. Best overall average value is highlighted in bold, while second best is underlined.
9 Additional Metrics
To study the calibration of our approach with respect to the state-of-the-art detectors as well as to facilitate comparison across popular metrics in the field, we expand the analysis of the main paper by computing the balanced accuracy on the $0.5$ threshold and the average precision metrics. We report the results across our test set of 5 sources of real images and 13 generative models in Tab. 8 and Tab. 9 respectively. As we see in the former, our approach, based on spectral learning, not only provides significant discriminative ability, but is also better calibrated around the common $0.5$ threshold, providing an absolute increase of 2.5% in balanced accuracy. Moreover, SPAI achieves an absolute increase of 3.5% in terms of average precision, reinstating its superior discriminative ability.
@@ 11 Qualitative Evaluation
We perform a qualitative evaluation of our approach across all the considered datasets and present samples for the 13 generative models in Figs. 12, 13, 14, 8, 9, 15, 16, 10, 17, 18, 19, 11 and 20 as well as for the 5 sources of real images in Figs. 21, 22, 23, 24 and 25. As we see, our approach accurately detects images generated by all the considered generative approaches, depicting a diverse set of topics and incorporating different levels of visual fidelity and aesthetics. At the same time, SPAI accurately classifies real images originating from all the employed sources. Therefore, employing spectral learning enables our architecture to not rely on some high-level semantics, but to effectively detect the subtle inconsistencies introduced by the generative models, to distinguish between AI-generated and real imagery.
12 Source Code
To facilitate the reproduction of our results as well as further research in the field we make publicly available our source code, data and trained models on https://mever-team.github.io/spai.